1. During the Pandemic a curious Science 104 student pitches a baseball into the air with an initial velocity of 24.5 m/s at an angle of 36.9° above the horizontal. He intends to find the ff:

a) Total time the baseball is in the air. Ans._______________

b) Range (Horizontal distance traveled) Ans._______________

c) Time when the baseball is at maximum height. Ans.¬¬¬¬¬¬_______________

d) Maximum height the baseball reaches. Ans._______________

e) Velocity of the baseball as it hits the ground. Ans._______________

Show Solution:


2. What angle will produce the farthest range? Justify your answers by showing a series of computations with angle θ0 as the variable summarizing your results by plotting horizontal range (R ) versus vertical; height (y). Hint V0 is constant and you can just assume any values)


Show Solгtion:


1
Expert's answer
2020-10-19T13:21:22-0400

1)

a)


"t=\frac{2(24.5)\sin{36.9}}{9.8}=3.00\ s"

b)


"R=\frac{(24.5)^2\sin{(2\cdot36.9)}}{9.8}=58.8\ m"

c)


"t'=0.5t=1.50\ s"

d)


"H=\frac{(24.5)^2\sin^2{36.9}}{2\cdot9.8}=11.0\ m"

e) Final velocity is 24.5 m/s at an angle of 36.9° below the horizontal.


2) 45° above the horizontal.